OpenStudy (sheetal):

find domain & range of y=x2+x-2/x2-x-2

OpenStudy (anonymous):

y=x2+x-2/x2-x-2 x2+x-2=x2+2x-x-2=(x-1)(x+2) y=(x-1)(x+2)/x2-x-2 if x=0 y=-1*2/-2=1 if y = 0 0=(x-1)(x+2) x=-2,1 range \[-2\le y \le1\] domain \[0\le x \le 1\]